# PT-Symmetry in $2\times 2$ Matrix Polynomials Formed by Pauli Matrices

### v1

@@ Abstract
Abstract: $2\times 2$ matrix polynomials of the form $P_{n}(z)=\Sigma^{n}_{j=0}\,\sigma_{j}\,z^{j}$, for the cases $n=1,2,3$ are constructed, and the nature of PT-symmetry is examined across different points $z=(x,y)$ in the complex plane. The PT-symmetric properties of $P_{n}(z)$ can be characterized by two functions, denoted by $s(x,y)$ and $h(x,y)$. If the trace of the matrix polynomial is real, then the points at which it can exhibit PT-symmetry are defined by the family of curves $s(x,y)=0$. Additionally, at points where the function $h(x,y)\geq 0$, the matrix polynomial exhibits unbroken PT-symmetry; otherwise, it exhibits broken PT-symmetry. The intersection points of the curves $s(x,y)=0$ and $h(x,y)=k$, for a given $k\in\mathbb{R}$, are shown to lie on an ellipse, hyperbola, two lines passing through the origin, or a straight line, depending on the nature of PT-symmetry of the matrix polynomial. The PT-symmetric behaviour of $P_{n}(z)$ at the zeros of the matrix polynomial is also studied.
@@ 1 Introduction
In the last part, we examine a trace-less matrix polynomial denoted by $Q_{10}(z)=z^{7}\,(P_{3}(z)-\sigma_{0})$ of degree $10$. By plotting the intersection points of the families of curves associated with $Q_{10}(z)$ defined by $s(x,y)=0$ and $h(x,y)=k$ for $k=-1,0,1$, we demonstrated that these intersection points lie on an ellipse. Similarly, for $k=\pm 0.0001$, some of the intersection points of the families of curves are shown to lie on an ellipse, while other points can be seen lie on a different ellipse. This result may suggest that intersection points are likely to lie on an ellipse when the absolute value of $k$ is sufficiently large for matrix polynomial of the form $z^{m}\,(P_{3}(z)-\sigma_{0})$ for any $m\geq 1$. Meanwhile, for a small value of $k$, the intersection points may lie on different ellipses.
@@ 2.4 PT-Symmetry in matrix polynomial of degree 10
Since $Q_{10}(z)$ is traceless, it is PT-symmetric at all points where $s(x,y)=0$. The family of curves $s(x,y)=0$ and $h(x,y)=k:k\neq 0$ intersect at 20 points and when $k=0$ intersection points are reduced to 5. The eigenvalues of the matrix polynomial at the intersection points are given by $E=\pm\sqrt{k}$ and $E=\pm i\sqrt{|k|}$, for $k>0$ and $K<0$, respectively. When $|k|\geq 1$, the intersection points lie on an ellipse as shown in Fig-12-(a) (for $k=-1$) and Fig-12-(c) (for $k=1$). When $k=0$, the curves intersect at the points given by $\{(x,y),(x,-y),(-x,y),(-x,-y),(0,0)\}$: where $x=\frac{1}{2}$ and $y=\frac{\sqrt{3}}{2}$ as shown in Fig-12-(b). Around the value of $|k|=0.5$, all the intersection points are observed to lie on a slightly deformed ellipse, as shown in Fig-13-(a)(for k= -0.5) and Fig-13-(b)(for k=0.5). As the value of $|k|$ approaches zero, the intersection points are observed to not lie on an ellipse; instead, 4 points lie on one ellipse, and the remaining 16 points are shown to lie on another ellipse, as shown in Fig-13-(c)(for k=-0.0001) and Fig-13-(d)(for k=0.0001). This structure continues until $k=0$, and when $k=0$, the 5 intersection points lie on two straight lines passing through the origin.
@@
This result may suggest that for matrix polynomial of the form $z^{m}\,(P_{3}(z)-\sigma_{0})$, all the intersection points are likely to lie on an ellipse when the absolute value of k is sufficiently large (i.e., not close to zero) for any $m\geq 1$. In contrast, for small values of k, the intersection points are likely to lie on different ellipses.
@@ 3 Summary
Matrix polynomial $\tilde{P}_{n}(z)$ are constructed by permuting the coefficient matrices of $P_{n}(z)$ for $n=1,2,3$ and are found to not possess PT-symmetry at the zeros of the matrix polynomials except for $n=1$. A matrix polynomial of degree 10 of the form $z^{m}\,(P_{3}(z)-\sigma_{0})$ is also examined, leading to the speculation that the points where it can have eigenvalues ±$\pm\sqrt{k}$ or ±$\pm i\sqrt{k}$ (with k positive and sufficiently large) lie on an ellipse for any $m\geq 1$. When the value of k is close to zero, the points may lie on different ellipses.
Acknowledgment

### v2

@@ Abstract
Abstract: $2\times 2$ matrix polynomials of the form $P_{n}(z)=\sum^{n}_{j=0}\,\sigma_{j}\,z^{j}$, for the cases $n=1,2,3$ are constructed, and the nature of PT-symmetry is examined across different points $z=(x,y)$ in the complex plane. The PT-symmetric properties of $P_{n}(z)$ can be characterized by two functions, denoted by $s(x,y)$ and $h(x,y)$. If the trace of the matrix polynomial is real, then the points at which it can exhibit PT-symmetry are defined by the family of curves $s(x,y)=0$. Additionally, at points where the function $h(x,y)\geq 0$, the matrix polynomial exhibits unbroken PT-symmetry; otherwise, it exhibits broken PT-symmetry. The intersection points of the curves $s(x,y)=0$ and $h(x,y)=k$, for a given $k\in\mathbb{R}$, are shown to lie on an ellipse, hyperbola, two lines passing through the origin, or a straight line, depending on the nature of PT-symmetry of the matrix polynomial. The PT-symmetric behaviour of $P_{n}(z)$ at the zeros of the matrix polynomial is also studied.
@@ 1 Introduction
[…]
@@ 2.4 PT-Symmetry in matrix polynomial of degree 10
Since $Q_{10}(z)$ is traceless, it is PT-symmetric at all points where $s(x,y)=0$. The family of curves $s(x,y)=0$ and $h(x,y)=k:k\neq 0$ intersect at 20 points and when $k=0$ intersection points are reduced to 5. The eigenvalues of the matrix polynomial at the intersection points are given by $E=\pm\sqrt{k}$ and $E=\pm i\sqrt{|k|}$, for $k>0$ and $K<0$, respectively. When $|k|\geq 1$, the intersection points lie on a slightly deformed ellipse as shown in Fig-12-(a) (for $k=-1$) and Fig-12-(c) (for $k=1$). When $k=0$, the curves intersect at the points given by $\{(x,y),(x,-y),(-x,y),(-x,-y),(0,0)\}$: where $x=\frac{1}{2}$ and $y=\frac{\sqrt{3}}{2}$ as shown in Fig-12-(b). Around the value of $|k|=0.5$, all the intersection points are observed to lie on a slightly deformed ellipse, as shown in Fig-13-(a)(for k= -0.5) and Fig-13-(b)(for k=0.5). As the value of $|k|$ approaches zero, the intersection points are observed to not lie on a single ellipse or a single deformed ellipse; instead, 4 points lie on one ellipse, and the remaining 16 points are shown to lie on another deformed ellipse, as shown in Fig-13-(c)(for k=-0.0001) and Fig-13-(d)(for k=0.0001). This structure continues until $k=0$, and when $k=0$, the five intersection points lie on two straight lines passing through the origin.
@@
This result may suggest that for matrix polynomial of the form $z^{m}\,(P_{3}(z)-\sigma_{0})$, all the intersection points are likely to lie on an ellipse or a slightly deformed ellipse when the absolute value of k is sufficiently large (i.e., not close to zero) for any $m\geq 1$. In contrast, for small values of k, the intersection points are likely to lie on different ellipses or deformed ellipses.
@@ 3 Summary
Matrix polynomial $\tilde{P}_{n}(z)$ are constructed by permuting the coefficient matrices of $P_{n}(z)$ for $n=1,2,3$ and are found to not possess PT-symmetry at the zeros of the matrix polynomials except for $n=1$. A matrix polynomial of degree 10 of the form $z^{m}\,(P_{3}(z)-\sigma_{0})$ is also examined, leading to the speculation that the points where it can have eigenvalues ±$\pm\sqrt{k}$ or ±$\pm i\sqrt{k}$ (with k positive and sufficiently large) lie on an ellipse or a slightly deformed ellipse for any $m\geq 1$. When the value of k is close to zero, the points may lie on different ellipses or deformed ellipses.
Acknowledgments